# ORION: Option-Regularized Deep Reinforcement Learning for Cooperative Multi-Agent Online Navigation

### v2

@@ IV-C Real-World Experiments
Finally, we deploy ORION on two ground vehicles (maximum speed $0.5\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$) equipped with Livox Mid-360 3D LiDARs for odometry and mapping, operating in a $30\text{\,}\mathrm{m}$$\times$$10\text{\,}\mathrm{m}$ office environment. We use OctoMap with an occupancy resolution of $0.4\text{\,}\mathrm{m}$ and a waypoint resolution of $0.5\text{\,}\mathrm{m}$ to construct the online map. As shown in Fig. 6, we evaluate ORION under two different start–target configurations that showcase distinct cooperative behaviors enabled by our framework. In both settings, the team travels approximately $20\text{\,}\mathrm{m}$ within $90\text{\,}\mathrm{s}$. In case 1, agent 1 (red) benefits from areas previously explored by agent 2 (yellow), demonstrating passive map sharing, where information gathered by one agent directly assists another without explicit coordination. In case 2, agent 1 exhibits post-arrival cooperation: after reaching its target, it moves toward the vicinity of target 2 to provide updated observations. Once target 2 enters the shared belief map, agent 2 immediately returns to its own target without redundant exploration, maintaining a compact travel distance while acquiring critical, up-to-date environmental information. Together, these real-world results highlight ORION’s ability to coordinate through both passive map sharing and active post-arrival assistance, underscoring its practicality for multi-robot deployment.
@@ V Conclusion
In this paper, we presented ORION, a deep reinforcement learning framework for cooperative multi-agent online navigation in partially changed environments. By fusing prior maps with online perception through a shared graph encoder and leveraging a novel option-critic structure, ORION enables agents to switch adaptively between target-directed navigation and cooperative assistance in a decentralized way. Our dual-stage cooperation strategy further enhances team-level performance by allowing arrived agents to contribute additional information that helps teammates reach their targets more efficiently. Extensive experiments on large-scale maze-like maps, high-fidelity Gazebo simulations, and physical robot deployments demonstrate that ORION achieves robust, real-time cooperation and consistently outperforms state-of-the-art baselines across varying team sizes.
Although ORION effectively handles dynamic map discrepancies, a current limitation lies in its assumption of reliable map sharing among agents. In future work, we aim to extend ORION to scenarios with limited or intermittent communication, where agents must reason about uncertain teammates’ beliefs and establish adaptive map-sharing and coordination strategies under bandwidth/latency constraints.

### v3

@@ IV-C Real-World Experiments
Finally, we deploy ORION on two ground vehicles (maximum speed $0.5\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$) equipped with Livox Mid-360 3D LiDARs for odometry and mapping, operating in a $30\text{\,}\mathrm{m}$$\times$$10\text{\,}\mathrm{m}$ office environment. We use OctoMap with an occupancy resolution of $0.4\text{\,}\mathrm{m}$ and a waypoint resolution of $0.5\text{\,}\mathrm{m}$ to construct the online map. During online execution, all robots share their locally built current maps, allowing the system to periodically merge them into a global current map. The unknown regions in this global current map are filled with prior-map information, producing the combined map used by all robots. As shown in Fig. 6, we evaluate ORION under two different start–target configurations that showcase distinct cooperative behaviors enabled by our framework. In both settings, the team travels approximately $20\text{\,}\mathrm{m}$ within $90\text{\,}\mathrm{s}$. In case 1, agent 1 (red) benefits from areas previously explored by agent 2 (yellow), demonstrating passive map sharing, where information gathered by one agent directly assists another without explicit coordination. In case 2, agent 1 exhibits post-arrival cooperation: after reaching its target, it moves toward the vicinity of target 2 to provide updated observations. Once target 2 enters the shared belief map, agent 2 immediately returns to its own target without redundant exploration, maintaining a compact travel distance while acquiring critical, up-to-date environmental information. Together, these real-world results highlight ORION’s ability to coordinate through both passive map sharing and active post-arrival assistance, underscoring its practicality for multi-robot deployment.
@@ V Conclusion
In this paper, we presented ORION, a deep reinforcement learning framework for cooperative multi-agent online navigation in partially changed environments. By fusing prior maps with online perception through a shared graph encoder and leveraging a novel option-critic structure, ORION enables agents to switch adaptively between target-directed navigation and cooperative assistance in a decentralized way. Our dual-stage cooperation strategy further enhances team-level performance by allowing arrived agents to contribute additional information that helps teammates reach their targets more efficiently. Extensive experiments on large-scale maze-like maps, high-fidelity Gazebo simulations, and physical robot deployments demonstrate that ORION achieves robust, real-time cooperation and consistently outperforms state-of-the-art baselines across varying team sizes. A current limitation lies in its assumption of reliable map sharing among agents. In future work, we plan to explicitly investigate more realistic settings with noisy communication/imperfect localization, where agents must reason about uncertain teammates’ beliefs and establish adaptive map-sharing and coordination strategies.